I need help solving this Log equation
4(1.02)^x=3(1.0)^x

Answer by lwsshak3(11628)

4(1.02)^x=3(1.0)^x=3 (1 raised to any number=1)
4(1.02)^x=3
(1.02)^x=3/4
xlog(1.02)=log3-log4
x=(log3-log4)/log(1.02)
x≈-14.5275